Charlotte (age 40) is a surviving spouse and provides all of the support of her four minor children, who live with her. Charlotte also maintains the household in which her parents live, and she furnished 60% of their support. Besides interest on City of Miami bonds in the amount of $5,500, Charlotte’s father received $2,400 from a part-time job. Charlotte earns an $80,000 salary, a short-term capital loss of $2,000, and a cash prize of $4,000 at a church raffle. Charlotte reports itemized deductions of $10,500.

a. Compute Charlotte's taxable income.

b. Using the Tax Rate Schedules, tax liability for Charlotte is $ _________ for 2018.

c. Compute Charlotte's child and dependent tax credit.

Basic Standard Deduction Amounts Filing Status 2018 Single $12,000 Married, filing jointly 24,000 Surviving spouse 24,000 18,000 Head of household Married, filing separately 12,000
2018 Tax Rate Schedules Single-Schedule X Head of household-Schedule Z If taxable of the If taxable of the income is: But not income is: But not amount amount The tax is: The tax is: Over Over over over over- over $ 9,525 $13,600 $ 10% 0 10% 0 $1,360.00+ 12% 9,525 38,700 13,600 51,800 13,600 952.5012% 9,525 82,500 4,453.5022% 51,800 82,500 38,700 38,700 5,944.0022% 51,800 14,089.50 24% 82,500 82,500 82,500 82,500 12,698.0024% 157,500 157,500 30,698.00 32% 157,500 157,500 157,500 157,500 200,000 32,089.5032% 200,000 200,000 500,000 200,000 200,000 500,000 200,000 45,689.50 35% 44,298.0035% 500,000 500,000 500,000 500,000 150,689.50+37% 149,298.00+37% . ....... Married filing jointly or Qualifying widow(er)- Schedule Y-1 Married filing separately-Schedule Y-2 f taxable of the If taxable of the income is: But not income is: But not amount amount The tax is: The tax is: Over Over over- Over- over over $ 19,050 $ 9,525 $ 10% ...10% S $ 19,050 38,700 $ 9,525 9,525 19,050 77,400 1,905.0012% 952.5012% 77,400 165,000 8,907.0022% 77,400 38,700 82,500 4,453.50 22% 38,700 28,179.00 24% 165,000 315,000 165,000 82,500 157,500 14,089.5024% 82,500 315,000 157,500 315,000 400,000 64,179,00+ 32% 157,500 200,000 32,089.5032% 91,379.00 35% 600,000 400,000 300,000 400,000 200.000 45,689.5035% 200,000 161,379.00 37% 600,000 600,000 300,000 300,000 80,689.5037%
Answer #1

Part A

Charlotte's AGI = salary – short term capital loss + cash prize = 80000-2000+4000 = $82000

Standard deduction for surviving spouse is $24000 and itemized deduction is $10500 (given), the greater amount from standard deduction and itemized deduction should be deducted from AGI to get taxable income. As the amount of standard deduction is greater, it will be deducted from AGI to get taxable income.

Taxable income = AGI – standard deduction = 82000 – 24000 = $58000

Part B

tax liability for Charlotte = $6579

using tax rate schedules for surviving spouse, tax liability = 1905+(12%*(58000-19050)) = $6579

Part C

Charlotte's child and dependent tax credit = (4*2000)+500 = $8500

Dear Student,

If any doubts, then feel free to ask.